y" - y = e2x[3 tan e+ 3 (sec e)^2]


1
Expert's answer
2021-04-15T07:20:02-0400



y"-y = e2x[3tan e+ "3sec e\over 2" ]

=Ce2x


"\implies" m2-1=0


(m+1)(m-1)=0

m= 1 , -1

complementary factor y

y= Acosh x + Bsinh x


The PI is y = De2x , y" = 4De2x


"\implies" 4De2x- De2x = Ce2x


"\implies" D="C\over3"


"\implies"


y= "e^{2x}[3tan e+ {3sece \over2} ]\over3"


We know

y = C.F + P.I


y= Acosh x + Bsinh x +

"e^{2x}[3tan e+ {3sece \over2} ]\over3"